A political discussion group consists of 5 Democrats and 7 Republicans. Four members are selected to attend a conference. Find the probability that the selected group will consist of all Republicans.
Group of answer choices
1
Expert's answer
2020-12-07T05:09:47-0500

"P(All Republicans)=\frac{C_7^4*C_5^0}{C_{12}^4}=\frac{35*1}{792}=\frac{355}{495}=0.0707."